OpenStudy (anonymous):

how do i Simplify (-4x)^4

OpenStudy (jiteshmeghwal9):



Parth (parthkohli):

A simpler way to think about it is \((-4x) \times (-4x) \times (-4x) \times (-4x)\).

Parth (parthkohli):

What is \((-4x) \times (-4x)\)?

OpenStudy (anonymous):

16x

Parth (parthkohli):

No! The \(x\)'s also are multiplied.

Parth (parthkohli):

\((-4x) \times (-4x) = -4 \times -4 \times x \times x\)

OpenStudy (anonymous):

huh

OpenStudy (jiteshmeghwal9):